OpenStudy (crashonce):

4cosx+3sinx=-1, solve for x

OpenStudy (shamim):

4cos x=-1-3sin x

OpenStudy (shamim):

Right?

OpenStudy (shamim):

After squaring both sides 16*cos^2x=1+6sin x+9sin^2x

OpenStudy (shamim):

We know Cos^2x=1-sin^2x

OpenStudy (shamim):

Use it

OpenStudy (shamim):

Response plz!!!